# AutoVP: An Automated Visual Prompting Framework and Benchmark

### v1

@@ 6 Limitations
This work is subject to some limitations. First, our optimization process did not include certain hyperparameters, notably learning rate and weight decay. This omission stemmed from our primary focus on identifying the best configurations for VP training, and because including such hyperparameters would have greatly increased execution workload. In addition, when it comes to choosing the best pre-trained model to fine-tune on the target dataset, You et al. (2021) also argued that, in general, the sophisticated fine-tuning techniques (e.g., regularization) would not change the ranking of pre-trained models in downstream tasks. Nonetheless, we conducted supplementary tuning experiments encompassing various learning rates and weight decays, the results of which can be found in Table 8 (within Appendix B.10). In practice, we suggest enabling the tuning of learning rates, weight decay, and other model-specific parameters after the initial hyperparameter tuning phase of AutoVP.
Another limitation pertains to the scope of AutoVP, which is oriented toward classification tasks. Recent research has demonstrated the success of visual prompts in generative tasks (Ramesh et al., 2021; 2022; Liu & Chilton, 2022; Bar et al., 2022; Sohn et al., 2023). However, expanding support for generative tasks will require accommodation of their distinctive requirements, e.g., via the integration of a suitable pre-trained generative model, such as GANs (Goodfellow et al., 2014), VAEs (Kingma & Welling, 2013), or diffusion models (Ho et al., 2020), along with tailored prompt design. Certainly, our results imply that there are many avenues for VP research that merit further exploration.
@@ 7 Conclusion
Appendix A Implementation Details
@@ A.2 Output Label Mappings of AutoVP
As mentioned in Section 3, AutoVP incorporates four output label mappings: frequency mapping (FreqMap), iterative mapping (IterMap), semantic mapping (SemanticMap), and fully connected layer mapping (FullyMap). In the following, we provide more details and the visual illustrations (Figure 6) of each mapping algorithm; we also illustrate the weight initialization specifically designed for the scenario where CLIP is used as the pre-trained model and FullyMap is utilized as the output mapping (more details can be found in Section 4.2).
@@
Figure 6: Output Label Mapping. Illustration of four output mapping methods: (I) Frequency Mapping, (II) Iterative Mapping, (III) Semantic Mapping, and (IV) Fully Connected Layer Mapping. (V) represents the weight initialization technique employed in the FullyMap when using CLIP as the pre-trained model.
In the FullyMap approach, we utilize a linear layer $L_{\text{t}}=w\cdot L_{\text{s}}+b$ with weights $w$ and bias $b$ to learn the mapping, enabling the transformation of the output source logits $L_{\text{s}}$ to target logits $L_{\text{t}}$. As for weight initialization (WI), it is employed in the case of CLIP with FullyMap. This technique involves setting the weights between the target labels and their default templates as 1, while setting the rest to 0. With WI, the linear layer can achieve a favorable initial mapping state, thereby expediting the process of obtaining a good mapping relation.
@@ A.3 AutoVP Tuning Process
In Figure 7, there are three stages involved in the tuning process, while the Visual Prompt component depicted in Figure 1 is not involved in the tuning process, as it does not contain any hyper-parameters. During the Input Scaling stage, the initial scale of the input image is determined, and users can choose whether to learn the resize scale during training. In the Pre-trained Classifier stage, users have the option to select from four pre-trained models to serve as the feature extractor. The Output Label Mapping stage offers four mapping methods to choose from. For FreqMap, IterMap, and SemanticMap, users can specify the number of source classes that are mapped to a single target class.
Figure 7: Hyper-Parameter Tuning Selection. Illustration of the end-to-end hyper-parameter tuning process in AutoVP with a total of 222 possible configurations.
Appendix B Experiments
In this section, we provide detailed experiments discussed in the main paper, covering several topics, including: 1) experimental settings (Section B.1 and B.2), 2) additional AutoVP experiments (Section B.3, B.4, and B.5), 3) methods comparisons (Section B.6), 4) pre-trained models comparisons (Section B.7, and B.8), 5) hyper-parameter tuning (Section B.9, B.10, and B.11), and 6) computing resources (Section B.12).
@@ B.1 The Twelve Downstream Datasets
To assess the efficacy of the proposed AutoVP, we selected the following twelve datasets for our experiments. The detailed information is shown in 3.
@@
Table 3: Dataset Setting
@@
SVHN 10 73257 26032 128
EuroSAT 10 13500 8100 128
Flowers102 102 4093 2463 64
CIFAR100 100 50000 10000 128
UCF101 101 7639 3783 128
DTD 47 2820 1692 32
FMoW 62 76863 22108 128
GTSRB 43 26640 12630 128
CIFAR10 10 50000 10000 128
Food101 101 75750 25250 128
OxfordIIITPet 37 3680 3669 128
ISIC 7 4990 555 128
@@ B.2 Baselines Details
B.3 Data Scalability
In the data scalability experiments, Figure 8 illustrates the performance of AutoVP and LP on each dataset under various data usage proportions. The corresponding settings can be found in Table 4. When the chosen pre-trained model is fixed to CLIP, AutoVP outperformed LP in scenarios with limited data on most of the datasets. In some datasets, such as SVHN, GTSRB, and EuroSAT, AutoVP performs better than LP for all proportions of data. Besides, in OxfordIIITPet, CIFAR100, DTD, Food101, UCF101, and Flowers102, AutoVP also obtains promising performance compared to LP on a small amount of data. This suggests that AutoVP will be a more suitable and effective solution than LP, especially when training data is limited.
Figure 8: Data Scalability. The charts present the accuracy of AutoVP and LP with varying percentages of data usage: 100%, 50%, 25%, 10%, and 1%.
Table 4: Data Scalability Settings. The table shows the settings for data scalability experiments with data usage ranging from 100% to 1%. The source pre-trained model is fixed to CLIP. The notation “Mapping-$m$” represents mapping $m$ source classes to each target class.
Dataset 100% 50% 25% 10% 1%
B.4 Downstream Dataset Analysis (ID/OOD vs. Accuracy Gain)
To further understand how the downstream dataset distribution influences the performance of visual prompting. We conduct experiments to observe the relation between accuracy gain and dataset characteristics. When using CLIP as the pre-trained model, we define confidence score, obtained by averaging the maximum softmax probability of predictions across the entire training set, as an indicator of dataset characteristics. For other vision pre-trained models, the FID score is used to measure the dissimilarity between the downstream dataset and the ImageNet-1K dataset, i.e. the degree of out-of-distribution (OOD). We present the Confidence/FID scores of each dataset in Figure 9, providing insights into their OOD characteristics. Furthermore, Figure 10 demonstrates the accuracy gain for each dataset when using ResNet18 as the pre-trained model. The experimental results show that when using AutoVP, datasets with a higher degree of OOD tend to benefit from more accuracy gains.
Figure 9: Confidence Score and FID Score. We sort the datasets by the degree of out-of-distribution (OOD), where the one closer to the left indicates higher similarity (in-distribution, ID) to the training data of the pre-trained model, while the one closer to the right indicates greater dissimilarity (OOD).
Figure 10: Accuracy Gains with Resnet18. The gains are calculated by taking the difference between the performance of AutoVP and LP or Non-VP scenario.
Additionally, Figure 11 showcases the testing accuracy on CLFAR10-C using the ResNet18 pre-trained model. Similarly, AutoVP exhibits superior performance/robustness compared to Linear Probing (LP).
Figure 11: Models robustness with ResNet18. The comparison of accuracy drop on the CIFAR10-C dataset between AutoVP and LP.
B.5 The Impact of Text Encoder in CLIP
Figure 12 illustrates the impact on accuracy when incorporating or excluding the CLIP text encoder. On average, this configuration results in a significant decrease in accuracy of approximately 12%, highlighting the crucial role of the text encoder in VP with CLIP.
Figure 12: The Accuracy Difference of AutoVP Variations. The chart shows the difference in accuracy between AutoVP variations with and without the CLIP text encoder. The numbers above the bars indicate the accuracy difference.
B.6 Comparison of AutoVP, Linear Probing, and Full Fine-Tuning
When comparing the performance of AutoVP to that of linear probing (LP) and full fine-tuning (FF), it’s crucial to acknowledge the significant discrepancy in terms of trainable parameter size. As FF involves a parameter size that is roughly 100 to 1000 times larger (as shown in Table 12). Nevertheless, it’s worth noting that the differences in performance between VP and FF are relatively minor in certain datasets. For instance, in the case of EuroSAT, both AutoVP and FF achieve a 96% accuracy (as demonstrated in Table 5). This observation suggests that VP retains its advantages even when faced with such substantial variations in parameter size.
Table 5: Comparison of AutoVP, Linear Probing (LP), and Full Fine-Tuning (FF). With the EuroSAT dataset and using CLIP as the pre-trained model, the table displays the testing accuracy, execution time, and trainable parameter size associated with each respective method.
Experimental Info. AutoVP LP FF
Accuracy (%) 96.84 94.62 96.78
Execution Time (second) 2448 2370 3081
Trainable Parameter Size (Million) 0.15 0.005 151.28
B.7 Fixed Pre-trained Model vs. Auto Pre-trained Model Selection
@@ B.6 Comparison of AutoVP, Linear Probing, and Full Fine-Tuning
We present additional experimental results to highlight the effectiveness of AutoVP. In addition to the comparisons with CLIP and VP baselines discussed in Section 4.1, we further evaluate the performance using ResNet18 as the pre-trained model and explore a scenario without any pre-trained model restrictions. The results using ResNet18 are presented in Table 6, while the results for the unrestricted scenario are provided in Table 7. These comparisons consistently demonstrate that AutoVP outperforms previous approaches, including LP and the state-of-the-art VP methods.
@@
Table 6: AutoVP with ResNet18. Comparison of VP test accuracy (%) using ResNet18 as the pre-trained model on 12 datasets.
@@
Table 7: AutoVP with source model selection. This table displays the best tuning setting without any restriction on the choice of pre-trained model, and shows the test accuracy (%) of AutoVP and the LP baseline of the chosen model across 12 datasets.
@@
B.8 Prompts in Frequency Domain
@@
We have analyzed the learned prompts using the best setting selected from AutoVP and have represented them in the frequency domain through Fast Fourier transformation (Brigham, 1988). In Figure 13(a), the prompting result of SVHN dataset with CLIP is notably distinct and achieves the highest testing accuracy among all the pre-trained models. In the frequency analysis (Figure 13(b)), the prompts are concentrated in the low-frequency domain (at the center of the plot), with CLIP displaying the most distinct structure and significantly larger magnitudes compared to the others. These results validate the efficient learning of prompts with CLIP, harnessing low-frequency features that generalize to the target domain.
@@
Figure 13: Prompting Analysis of SVHN Using Various Pre-trained Models. (a) Frame-shape prompts learned with the best settings selected from AutoVP for a given pre-trained model. (b) Prompts in the frequency domain by Fast Fourier transformation.
B.9 The Preferences in Hyper-Parameter Tuning Selection
Figure 14 illustrates the average tuning results for all datasets across different settings, including image scale, mapping methods, pre-trained models, and so on (see Figure 7). In Figure 14(a), among the vision models (ResNet18, ResNeXt-IG, Swin-T), the FullyMap demonstrates superior performance compared to the other methods. However, for CLIP, a vision-language model, the FullyMap only shows a slight advantage over the others. Furthermore, Figure 14(b) indicates that larger initial scales, such as 1.0 or 1.5 (yielding square images with a width of $128\times 1.0$ or $128\times 1.5$), generally lead to better results when using FreqMap, IterMap, or FullyMap. Last, since both FreqMap and IterMap can configure the number ($m$) of source labels mapped to each target class, we found that increasing this count generally improves accuracy for the three vision models (see Figure 14(c)). However, for CLIP, mapping five source labels appears to be the optimal choice based on the average tuning results.
@@ B.9 The Preferences in Hyper-Parameter Tuning Selection
Figure 14: Average Tuning Results. The three charts display the average few-epoch tuning accuracy for different selected conditions.
B.10 Exploring Additional Tuning Axes

### v2

@@ 6 Limitations
This work is subject to some limitations. First, our optimization process did not include certain hyperparameters, notably learning rate and weight decay. This omission stemmed from our primary focus on identifying the best configurations for VP training, and because including such hyperparameters would have greatly increased execution workload. In addition, when it comes to choosing the best pre-trained model to fine-tune on the target dataset, You et al. (2021) also argued that, in general, the sophisticated fine-tuning techniques (e.g., regularization) would not change the ranking of pre-trained models in downstream tasks. Nonetheless, we conducted supplementary tuning experiments encompassing various learning rates and weight decays, the results of which can be found in Table E.3 (within Appendix E.3). In practice, we suggest enabling the tuning of learning rates, weight decay, and other model-specific parameters after the initial hyperparameter tuning phase of AutoVP. The tuning of fundamental hyperparameters could potentially be accelerated with recent advancements in utilizing generalization metrics to identify optimal hyperparameter configurations (Zhou et al., 2023a; b), a subject to be explored in future research.
Another limitation pertains to the scope of AutoVP, which is oriented toward classification tasks. However, we have extended its application to segmentation and detection tasks, as detailed in Appendix E.2. Furthermore, recent studies have demonstrated the success of visual prompts in generative tasks (Ramesh et al., 2021; 2022; Liu & Chilton, 2022; Bar et al., 2022; Sohn et al., 2023). Nevertheless, expanding support for generative tasks will require accommodation of their distinctive requirements, e.g., via the integration of a suitable pre-trained generative model, such as GANs (Goodfellow et al., 2014), VAEs (Kingma & Welling, 2013), or diffusion models (Ho et al., 2020), along with tailored prompt design. Certainly, our results imply that there are many avenues for VP research that merit further exploration.
@@ 7 Conclusion
Acknowledgments
The research described in this paper was conducted in the JC STEM Lab of Intelligent Design Automation, which is funded by The Hong Kong Jockey Club Charities Trust.
@@ Acknowledgments
[…]
Appendix A Implementation Details of AutoVP
@@ A.2 Output Label Mappings of AutoVP
As mentioned in Section 3, AutoVP incorporates four output label mappings: frequency mapping (FreqMap), iterative mapping (IterMap), semantic mapping (SemanticMap), and fully connected layer mapping (FullyMap). In the following, we provide more details of each mapping algorithm.
Figure A.1: Output Label Mapping. Illustration of four output mapping methods.
A.2.1 FreqMap (Frequency Mapping)
@@ 3 $\bf{M}\leftarrow\bf{0}^{K_{\text{s}}\times K_{\text{t}}}$     # mapping matrix
A.2.2 IterMap (Iterative Mapping)
A.2.3 SemanticMap (Semantic Mapping)
A.2.4 FullyMap (Fully Connected Layer Mapping)
FullyMap utilizes a linear layer $L_{\text{t}}=w\cdot L_{\text{s}}+b$ with weights $w$ and bias $b$ to learn the mapping, enabling the transformation of the output source logits $L_{\text{s}}$ to target logits $L_{\text{t}}$. As for weight initialization (WI), it is employed in the case of CLIP with FullyMap. This technique involves setting the weights between the target labels and their default templates as 1, while setting the rest to 0. With WI, the linear layer can achieve a favorable initial mapping state, thereby expediting the process of obtaining a good mapping relation.
@@ A.3 AutoVP Tuning Process
In Figure A.2, there are three stages involved in the tuning process, while the Visual Prompt component depicted in Figure 1 is not involved in the tuning process, as it does not contain any hyper-parameters. During the Input Scaling stage, the initial scale of the input image is determined, and users can choose whether to learn the resize scale during training. In the Pre-trained Classifier stage, users have the option to select from four pre-trained models to serve as the feature extractor. The Output Label Mapping stage offers four mapping methods to choose from. For FreqMap, IterMap, and SemanticMap, users can specify the number of source classes that are mapped to a single target class.
Figure A.2: Hyper-Parameter Tuning Selection. Illustration of the end-to-end hyper-parameter tuning process in AutoVP with a total of 222 possible configurations.
Appendix B Datasets and Baselines
@@ B.1 The Twelve Downstream Datasets
To assess the efficacy of the proposed AutoVP, we selected the following twelve datasets for our experiments. The detailed information is shown in B.1.
@@
Table B.1: Dataset Setting
@@
SVHN 10 73,257 26,032 128
EuroSAT 10 13,500 8,100 128
Flowers102 102 4,093 2,463 64
CIFAR100 100 50,000 10,000 128
UCF101 101 7,639 3,783 128
DTD 47 2,820 1,692 32
FMoW 62 76,863 22,108 128
GTSRB 43 26,640 12,630 128
CIFAR10 10 50,000 10,000 128
Food101 101 75,750 25,250 128
OxfordIIITPet 37 3,680 3,669 128
ISIC 7 4,990 555 128
@@ B.2 Baselines Details
Appendix C Additional Experiments with AutoVP
C.1 Fixed Pre-trained Model vs. Auto Pre-trained Model Selection
@@ Appendix C Additional Experiments with AutoVP
We present additional experimental results to highlight the effectiveness of AutoVP. In addition to the comparisons with CLIP and VP baselines discussed in Section 4.1, we further evaluate the performance using ResNet18 as the pre-trained model and explore a scenario without any pre-trained model restrictions. The results using ResNet18 are presented in Table C.1, while the results for the unrestricted scenario are provided in Table C.2. These comparisons consistently demonstrate that AutoVP outperforms previous approaches, including LP and the state-of-the-art VP methods.
@@
Table C.1: AutoVP with ResNet18. Comparison of VP test accuracy (%) using ResNet18 as the pre-trained model on 12 datasets.
@@
Table C.2: AutoVP with source model selection. This table displays the best tuning setting without any restriction on the choice of pre-trained model, and shows the test accuracy (%) of AutoVP and the LP baseline of the chosen model across 12 datasets.
@@
C.2 Data Scalability
In the data scalability experiments, Figure C.1 illustrates the performance of AutoVP and LP on each dataset under various data usage proportions. The corresponding settings can be found in Table C.3. When the chosen pre-trained model is fixed to CLIP, AutoVP outperformed LP in scenarios with limited data on most of the datasets. In some datasets, such as SVHN, GTSRB, and EuroSAT, AutoVP performs better than LP for all proportions of data. Besides, in OxfordIIITPet, CIFAR100, DTD, Food101, UCF101, and Flowers102, AutoVP also obtains promising performance compared to LP on a small amount of data. This suggests that AutoVP will be a more suitable and effective solution than LP, especially when training data is limited.
Figure C.1: Data Scalability. The charts present the accuracy of AutoVP and LP with varying percentages of data usage: 100%, 50%, 25%, 10%, and 1%.
Table C.3: Data Scalability Settings. The table shows the settings for data scalability experiments with data usage ranging from 100% to 1%. The source pre-trained model is fixed to CLIP. The notation “Mapping-$m$” represents mapping $m$ source classes to each target class.
Dataset 100% 50% 25% 10% 1%
C.3 Downstream Dataset Analysis (ID/OOD vs. Accuracy Gain)
To further understand how the downstream dataset distribution influences the performance of visual prompting. We conduct experiments to observe the relation between accuracy gain and dataset characteristics. When using CLIP as the pre-trained model, we define confidence score, obtained by averaging the maximum softmax probability of predictions across the entire training set, as an indicator of dataset characteristics. For other vision pre-trained models, the FID score is used to measure the dissimilarity between the downstream dataset and the ImageNet-1K dataset, i.e. the degree of out-of-distribution (OOD). We present the Confidence/FID scores of each dataset in Figure C.2, providing insights into their OOD characteristics. Furthermore, Figure C.3 demonstrates the accuracy gain for each dataset when using ResNet18 as the pre-trained model. The experimental results show that when using AutoVP, datasets with a higher degree of OOD tend to benefit from more accuracy gains.
Figure C.2: Confidence Score and FID Score. We sort the datasets by the degree of out-of-distribution (OOD), where the one closer to the left indicates higher similarity (in-distribution, ID) to the training data of the pre-trained model, while the one closer to the right indicates greater dissimilarity (OOD).
Figure C.3: Accuracy Gains with Resnet18. The gains are calculated by taking the difference between the performance of AutoVP and LP or Non-VP scenario.
Table C.4: Performance on Additional Datasets. The testing accuracy (%) for four additional datasets commonly found in VP research. The figure on the right side shows the accuracy gains.
Dataset AutoVP Setting AutoVP LP ILM CLIP-VP
SUN397 FullyMap, $p=16$ 65.4 70.9 61.2 60.5
StanfordCar FullyMap, $p=16$ 61.8 77.6 57.6 56.2
RESISC FullyMap, $p=17$ 88.5 91.7 86.6 84.5
CLEVR FullyMap, $p=16$ 83.0 68.0 83.1 81.4
Table 2 has encompassed the prevalent 12 datasets in VP research. Furthermore, to enable a more comprehensive comparison within a wider spectrum of VP research, we have included additional datasets—SUN397(Xiao et al., 2010), RESISC(Cheng et al., 2017), CLEVR(Johnson et al., 2017), and StanfordCar(Krause et al., 2013)—with their respective results available in Table C.4. Our findings consistently demonstrate AutoVP’s superior performance compared to ILM and CLIP-VP across the most of datasets. Especially notable is the substantial 15% increase in accuracy observed in the out-of-distribution (OOD) dataset, CLEVR, when compared to linear probing (LP). However, the in-distribution (ID) datasets, SUN397 and StanfordCar, showcased lower performance than LP, aligning with the trend of accuracy gain illustrated in Figure 5 in Section 5.
Appendix D Analysis of AutoVP Results
D.1 Prompts in Frequency Domain
@@
We have analyzed the learned prompts using the best setting selected from AutoVP and have represented them in the frequency domain through Fast Fourier transformation (Brigham, 1988). In Figure D.1(a), the prompting result of SVHN dataset with CLIP is notably distinct and achieves the highest testing accuracy among all the pre-trained models. In the frequency analysis (Figure D.1(b)), the prompts are concentrated in the low-frequency domain (at the center of the plot), with CLIP displaying the most distinct structure and significantly larger magnitudes compared to the others. These results validate the efficient learning of prompts with CLIP, harnessing low-frequency features that generalize to the target domain.
@@
Figure D.1: Prompting Analysis of SVHN Using Various Pre-trained Models. (a) Frame-shape prompts learned with the best settings selected from AutoVP for a given pre-trained model. (b) Prompts in the frequency domain by Fast Fourier transformation.
D.2 Output Mapping Analysis
Figure D.2: The Weight Distribution of FullyMap in DTD. The top 5 source labels exhibiting the highest weight values within the FullyMap pertain to (a) the label bumpy and (b) the label scaly.
Figure D.2 illustrates that FullyMap can be interpreted as a weighted combination of multiple source labels, where some human-readable features may exhibit similarity. For instance, in Figure D.2(a), bumpy shows similarities with custard apple, disk brake, and pineapple, while in Figure D.2(b), scaly shares similar features with boa constrictor, coho, and common iguana.
Figure D.3: The Correspondence Between Output Mapping Labels in DTD. Each column represents the mapping between the target label and its respective top-1 source label in FullyMap. (a) FullyMap (top-1 class having the largest weight) with Swin-T (second row). (b) IterMap-1 with Swin-T (third row).
Furthermore, when comparing FullyMap and IterMap, a significant accuracy gap is observed, with FullyMap achieving 69.96% and IterMap-1 only achieving 40.77%. However, in Figure D.3, IterMap has mapped to some classes that are indeed very close to the target. For instance, in Figure D.3(b), braided maps to knot, bubbly maps to bubble, and cobwebbed maps to spider web. This demonstrates that a mere combination of source labels is insufficient for achieving better performance; the weighting in the combination plays a crucial role, which is precisely what FullyMap accomplishes.
D.3 The Preferences in Hyper-Parameter Tuning Selection
@@
Figure D.4: Average Tuning Results. The three charts display the average few-epoch tuning accuracy for different selected conditions.
Figure D.4 illustrates the average tuning results for all datasets across different settings, including image scale, mapping methods, pre-trained models, and so on (see Figure A.2). In Figure D.4(a), among the vision models (ResNet18, ResNeXt-IG, Swin-T), the FullyMap demonstrates superior performance compared to the other methods. However, for CLIP, a vision-language model, the FullyMap only shows a slight advantage over the others. Furthermore, Figure D.4(b) indicates that larger initial scales, such as 1.0 or 1.5 (yielding square images with a width of $128\times 1.0$ or $128\times 1.5$), generally lead to better results when using FreqMap, IterMap, or FullyMap. Last, since both FreqMap and IterMap can configure the number ($m$) of source labels mapped to each target class, we found that increasing this count generally improves accuracy for the three vision models (see Figure D.4(c)). However, for CLIP, mapping five source labels appears to be the optimal choice based on the average tuning results.
Appendix E Ablation Studies
E.1 The Impact of Text Encoder in CLIP
Figure E.1 illustrates the impact on accuracy when incorporating or excluding the CLIP text encoder. On average, this configuration results in a significant decrease in accuracy of approximately 12%, highlighting the crucial role of the text encoder in VP with CLIP.
Figure E.1: The Accuracy Difference of AutoVP Variations. The chart shows the difference in accuracy between AutoVP variations with and without the CLIP text encoder. The numbers above the bars indicate the accuracy difference.
E.2 Visual Prompting in Segmentation and Detection Tasks
Figure E.2: LP and AutoVP in Segmentation and Detection Tasks.
Although AutoVP primarily focuses on classification tasks, we aimed to delve deeper into various vision tasks by incorporating models designed for different purposes. For segmentation tasks, we employ DeepLabV3 (Chen et al., 2017) as the pre-trained backbone, and for detection tasks, we utilize OWL-ViT (Minderer et al., 2022). To evaluate the capability of both in-distribution (ID) and out-of-distribution (OOD) datasets, we chose ISIC as the OOD dataset for both tasks. For ID datasets, we use Pets in the segmentation task and VOC (Everingham et al., 2015) in the detection task.
Our VP segmentation framework, depicted in Figure E.2 (a), integrates a FullyMap after the pre-trained model to facilitate pixel-wise classification using a custom class number. In contrast, the linear probing approach modifies the final 2D convolutional layer, and the results are depicted in Table E.1.
Table E.1: Performance on Segmentation Datasets. The IoU and pixel accuracy (%) of linear probing (LP) and AutoVP.
Dataset LP AutoVP
Pets IoU : 0.83, Pixel : 90.7% IoU : 0.77, Pixel : 86.9%
ISIC IoU : 0.64, Pixel : 78.1% IoU : 0.81, Pixel : 89.5%
We evaluated segmentation performance using two metrics: IoU (Intersection over Union) score and pixel-wise classification accuracy. AutoVP exhibited superior performance on both metrics on the ISIC dataset. Additionally, segmentation examples highlighted that predictions align more accurately with the ground truth mask when the prompt space is larger (see Figure E.3). However, in the ID dataset (Pets), VP performance was inferior to LP. This aligns with our findings in the classification task, where OOD datasets derived greater benefits from visual prompts.
Figure E.3: ISIC Segmentation. Performance of LP and AutoVP with different prompt sizes ($p$). It’s important to note that the performance calculation involves only the target image itself without the prompts, ensuring a fair comparison to LP. Specifically, the predicted mask is initially cropped to the region of the original target image and then resized to match the dimensions used in LP.
In the case of the detection task, as illustrated in Figure E.2 (b), the LP approach fine-tunes on the box predictor head, whereas the VP method incorporates trainable visual prompts. To achieve consistent box output, we have omitted the inclusion of output mapping. The results are shown in Table E.2. Similar to segmentation, the OOD dataset, ISIC, exhibits better performance than LP with a larger prompt size.
Table E.2: Performance on Detection Datasets. The IoU scores of linear probing (LP) and AutoVP.
Dataset Zero-shot Finetune box head VP
VOC IoU: 0.75 IoU: 0.76 IoU: 0.73
ISIC IoU: 0.52 IoU: 0.70 IoU: 0.80
E.3 Exploring Additional Tuning Axes